OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log 1/5 x. Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts. (4 points)

OpenStudy (anonymous):

\( \log \frac{1}{5}x\) or \(\log\frac{1}{5x}\)